# Track2Act: Predicting Point Tracks from Internet Videos enables Diverse Zero-shot Robot Manipulation

### v1

@@ 4 Experiment Setup
We focus our experiments for Track2Act  on in-the-wild manipulation scenarios where a mobile manipulator needs to manipulate objects in different living rooms, offices, and kitchens based on specified goals. For all the robot experiments, we use a Boston Dynamics Spot robot equipped with a manipulator (hand) and a front facing Intel RealSense camera [70]. We manipulate the arm through end-effector control based on the outputs of our policy.
@@ 4.1 Evaluation Details
Robot Experiments. As is the convention in goal-conditioned robot learning, we perform evaluations by quantifying success rate, where a successful trajectory is defined to be one where the final pose of the object in the scene to be manipulated is identical to the pose of the object in the goal image. We categorize results based on different levels of generalization, the definitions of which are inspired by prior works [1, 9, 2, 3]:
@@ 4.2 Baselines and Comparisons
Robot Experiments. We perform several comparisons with baselines and ablation studies for goal-conditioned robot manipulation. For baselines, we use the same embodiment-specific demonstrations as Track2Act, the goal-conditioned policy that predicts residuals over open-loop actions at each time-step (Algorithm 2).
@@
Track2Act(Open Loop) is the approach for track prediction followed by open-loop execution as described in Algorithm 1. This does not use any embodiment-specific data for training. To understand the benefit of predicting residuals over actions as opposed to predicting complete actions, we compare with an ablated variant Track2Act(actions; not residuals) that predicts actions $\hat{\mathbf{a}}_{t}$ directly without predicting residuals $\Delta\mathbf{a}_{t}$ and not relying on an open-loop plan as input.
@@ 4.3 Training Data
For training the track prediction model, we leverage diverse passive videos available on the web that are not collected by us. Specifically, we use human video clips from EpicKitchens [17] (clipping videos to ensure they are of 4-5 seconds duration), human videos on YouTube sourced in SmthSmthv2 [12], and large-scale robot videos released in RT1 data [2] and BridgeData [74]. To obtain ground-truth tracks for training the prediction model, we run Co-Tracker [11] on the resulting 400,000 video clips. Note that the robot datasets (RT1 and Bridge) are on completely different robots and scenarios than the robot we use for experiments (Spot). For training the residual policy, the embodiment-specific data we collect consists of $\sim 400$ trajectories obtained by tele-operating the Spot, for solving 10 tasks of manipulating everyday objects like doors, drawers, bottles, jugs. Note that this embodiment-specific data we collect is 3-4 orders of magnitude less than that what related works [2, 3, 1] require for policy learning.
@@ 5.1 Point Track Prediction Results
Figure 5: We show qualitative results of the track predictions for Track2Act on unseen initial and goal images across diverse datasets. Given specified points on the initial image we predict future tracks of these points, corresponding to the goal image. We can see that the predictions are plausible and correspond to manipulating the object(s) in the scene.
@@ 5.2 Robot Manipulation Results
Figure 7: Qualitative results showing robot executions (from a third person view) with the residual policy for different tasks with respect to the generalization levels defined in section 4.1. We show the first and last images of a rollout. The robot executions are best viewed as videos in the linked website.
@@
Track2Act(Open-Loop) 35% 25% 30% 25%
Track2Act(Ablation; actions not residuals) 70% 45% 30% 30%
Track2Act 70% 60% 55% 40%
@@ 5.3 Analysis of Failures
Here we discuss the failures displayed by Track2Act. For the open-loop plan based on residual transforms, the main failure modes we observe are inability to grasp the object at the right location, and inability to recover from intermediate failures. The residual policy corrects for these behaviors by virtue of leveraging some embodiment-specific data, and thus has higher success rates. We note that the success rate for higher levels of generalization CG and TG) is still not very high since these are very challenging settings and the residual policy sometimes fails by incorrectly grasping the object, getting stuck during the execution by trying to execute a non-feasible motion, or by executing a trajectory that does not conform with the goal image specified.
@@ 6 Discussion and Conclusion
In this paper, we developed Track2Act  - a framework for generalizable zero-shot robot manipulation by leveraging large-scale web video data to learn embodiment agnostic plans of how objects should be manipulated in a scene to satisfy a goal. We combined this with a small amount of embodiment-specific data to learn residual corrections over the predicted plans through a closed-loop policy. Our real world manipulation results across a range of diverse tasks with varying levels of generalization demonstrate the potential of scalably leveraging web data to predict plans for object manipulation. While our framework allows for strong generalization to unseen tasks in-the-wild, the tasks are still of short-horizon and involve manipulating a single object in the scene. It would be an interesting direction of future work to extend our framework for tackling long-horizon tasks that involve successive manipulations of multiple objects in the scene.
@@ 6.7 Details on baselines
We perform several comparisons with baselines and ablation studies for goal-conditioned robot manipulation. For baselines, we use the same embodiment-specific demonstrations as Ours, the goal-conditioned policy that predicts residuals over open-loop actions at each time-step (Algorithm 2).

### v2

@@ 4 Experiment Setup
We focus our experiments on in-the-wild manipulation scenarios where a mobile manipulator needs to manipulate objects in different living rooms, offices, and kitchens based on specified goals. For all the robot experiments, we use a Boston Dynamics Spot robot equipped with a manipulator (hand) and a front facing Intel RealSense camera [32]. We manipulate the arm through end-effector control based on the outputs of our policy.
@@ 4.1 Evaluation Details
Track Prediction. As is the convention in goal-conditioned robot learning, we perform evaluations by quantifying success rate, where a successful trajectory is defined to be one where the final pose of the object in the scene to be manipulated is identical to the pose of the object in the goal image. We categorize results based on different levels of generalization, the definitions of which are inspired by prior works [5, 4, 8, 73]:
@@ 4.2 Baselines and Comparisons
Robot Experiments. We perform several comparisons with baselines and ablation studies for goal-conditioned robot manipulation. For baselines, we use the same embodiment-specific demonstrations as Ours, the goal-conditioned policy that predicts residuals over open-loop actions at each time-step.
@@ 4.3 Training Data
For training the track prediction model, we leverage diverse passive videos available on the web that are not collected by us. Specifically, we use human video clips from EpicKitchens [10] (clipping videos to ensure they are of 4-5 seconds duration), and large-scale robot videos released in RT1 data [8] and BridgeData [61]. To obtain ground-truth tracks for training the prediction model, we run Co-Tracker [29] on the resulting 400,000 video clips. Note that the robot datasets (RT1 and Bridge) are on completely different robots and scenarios than the robot we use for experiments (Spot). For training the residual policy, the embodiment-specific data we collect consists of $\sim 400$ trajectories obtained by tele-operating the Spot, for solving 10 tasks of manipulating everyday objects like doors, drawers, bottles, jugs. Note that this embodiment-specific data we collect is 3-4 orders of magnitude less than that what related works [8, 73, 5] require for policy learning.
@@ 5.1 Point Track Prediction Results
Figure 5: We show qualitative results of the track predictions on unseen initial and goal images across diverse datasets. Given specified points on the initial image we predict future tracks of these points, corresponding to the goal image. We can see that the predictions are plausible and correspond to manipulating the object(s) in the scene.
@@ 5.2 Robot Manipulation Results
Figure 7: Qualitative results showing robot executions (from a third person view) with the residual policy for different tasks with respect to the generalization levels defined in section 4.1. We show the first and last images of a rollout. The robot executions are best viewed as videos in the supplementary zip.
@@
Ours (Open-Loop) 35% 25% 30% 25%
Ours (Ablation; actions not residuals) 70% 45% 30% 30%
Ours 70% 60% 55% 40%
@@ 5.3 Analysis of Failures
Here we discuss the failures displayed by our framework. For the open-loop plan based on residual transforms, the main failure modes we observe are inability to grasp the object at the right location, and inability to recover from intermediate failures. The residual policy corrects for these behaviors by virtue of leveraging some embodiment-specific data, and thus has higher success rates. We note that the success rate for higher levels of generalization CG and TG) is still not very high since these are very challenging settings and the residual policy sometimes fails by incorrectly grasping the object, getting stuck during the execution by trying to execute a non-feasible motion, or by executing a trajectory that does not conform with the goal image specified.
@@ 6 Discussion and Conclusion
In this paper, we developed a framework for generalizable robot manipulation by leveraging large-scale web video data to learn embodiment agnostic plans of how objects should be manipulated in a scene to satisfy a goal. We combined this with a small amount of embodiment-specific data to learn residual corrections over the predicted plans through a closed-loop policy. Our real world manipulation results across a range of diverse tasks with varying levels of generalization demonstrate the potential of scalably leveraging web data to predict plans for object manipulation. While our framework allows for strong generalization to unseen tasks in-the-wild, the tasks are still of short-horizon and involve manipulating a single object in the scene. It would be an interesting direction of future work to extend our framework for tackling long-horizon tasks that involve successive manipulations of multiple objects in the scene.
@@ 6.7 Details on baselines
We perform several comparisons with baselines and ablation studies for goal-conditioned robot manipulation. For baselines, we use the same embodiment-specific demonstrations as Ours, the goal-conditioned policy that predicts residuals over open-loop actions at each time-step (Algorithm LABEL:alg:residual).